From a box containing 4 black balls, 2 green balls, and 1 white ball, 3 balls is drawn in succession, each ball is not replaced in the box before the next draw is made. If X is the number of green balls drawn.

(a) What are the values for the set of x?
(b) What is the probability of having 2 green balls? (Fraction form)
(c) Make a probability distribution table.
1
Expert's answer
2020-12-10T14:19:54-0500

Well let`s explain my solution:

a) It is clear that x has interval [0,2]. if you draw three balls, then there may be not be a green ball or may have maximum 2 balls.

b) In a box we have 7 balls and 2 of them are green. if we divide number green balls to total number of balls, we will find probablity of that at least one green ball was drawn: I meant this--> "\frac{2}{7}" and for the second green ball: "\frac{1}{6}" because first green ball took out. If we multiply both of them, we will find the probability of having 2 green balls: "\frac{2}{7}" "\times" "\frac{1}{6}" = "\frac{1}{21}" --> this is answer.

c) Let`s construct a probability distribution table.

x | white | green | black

P(X=x) | "\frac{1}{7}" | "\frac{2}{7}" | "\frac{4}{7}" ---> this is distribution table.